# iNews: A Multimodal Dataset for Modeling Personalized Affective Responses to News

### v1

@@ Abstract
Current approaches to emotion detection often overlook the inherent subjectivity of affective experiences, instead relying on aggregated labels that mask individual variations in emotional responses. We introduce iNews, a novel large-scale dataset explicitly capturing subjective affective responses to news headlines. Our dataset comprises annotations from 291 demographically diverse UK participants across 2,899 multimodal Facebook news posts from major UK outlets, with an average of 5.18 annotators per sample. For each post, annotators provide multifaceted labels including valence, arousal, dominance, discrete emotions, content relevance judgments, sharing likelihood, and modality importance ratings (text, image, or both). Furthermore, we collect comprehensive annotator persona information covering demographics, personality, media trust, and consumption patterns, which explain 15.2% of annotation variance - higher than existing NLP datasets. Incorporating this information yields a 7% accuracy gain in zero-shot prediction and remains beneficial even with 32-shot. iNews will enhance research in LLM personalization, subjectivity, affective computing, and individual-level behavior simulation [^1] The dataset will be made publicly available at [HuggingFace Link]..
@@
Tiancheng Hu University of Cambridge th656@cam.ac.uk                        Nigel Collier University of Cambridge nhc30@cam.ac.uk
@@ 1 Introduction
Emotion is a quintessential example of the human subjectivity that permeates human experience, profoundly shaping our perceptions, decisions, and interactions (Zadra and Clore, 2011; Lerner et al., 2015, inter alia). As large language models (LLMs) achieve widespread global adoption, understanding and modeling these subjective affective responses becomes crucial for developing systems that can effectively personalize interactions and ensure alignment with diverse human values and preferences. Affective responses vary significantly across individuals, shaped by a complex interplay of individual and group-level factors, including age, gender, personality, cultural backgrounds, and lived experiences Kring and Gordon (1998); Costa and McCrae (2008); Charles and Carstensen (2010); Mesquita and Frijda (1992).
Despite the well-established understanding in psychology of the inherent subjectivity of emotion, many natural language processing (NLP) approaches to emotion detection fall short of embracing a user-centric perspective that accounts for individual differences. They often rely on aggregation techniques, such as majority voting or averaging, which obscure nuanced individual reactions and neglect to explicitly define the intended perspective of the analysis (e.g., writer vs. reader) Ovesdotter Alm (2011); Plank (2022); Cabitza et al. (2023). Such aggregation, without acknowledging individual variability and potential biases, risks generating misleading conclusions and hindering the development of personalized and user-centric systems.
@@
To address these limitations, we introduce iNews, a novel large-scale dataset specifically designed to capture the inherent subjectivity of affective responses to real-world news content (overview in Figure 1). Our dataset comprises fine-grained affective responses from 291 annotators to 2,899 Facebook posts from leading UK media outlets. The annotations include: valence-arousal-dominance (VAD) ratings Mehrabian and Russell (1974); Bradley and Lang (1994), Plutchik’s eight basic emotions Plutchik (1980), perceived post relevance, modality importance, and sharing likelihood. In addition, we collect a comprehensive set of annotator characteristics [^2] Throughout this work, we also use the term persona information to refer to the same concept. (e.g. demographics, personality, media consumption habits), drawing upon insights from the differential media effects literature.
@@
Our regression analysis confirms that annotator characteristics explain a substantial portion of the annotation variance (15.2% - higher than observed in any NLP dataset to date), highlighting the importance of incorporating individual differences when modeling subjective phenomena like affect. Furthermore, through an open-ended questionnaire with a subset of annotators ($N=20$), we identify nuanced patterns in how individuals experience and articulate their emotional reactions to news content, beyond the scope of our structured annotations and survey.
@@
In a case study demonstrating the practical value of this rich persona information, we show that incorporating annotator characteristics can improve LLM predictions of individual-level affective responses by up to 7% in accuracy in zero-shot settings, although overall accuracy remains relatively modest (around 40%). When comparing different input modalities (image vs. text), we find that image inputs typically outperform text in zero-shot scenarios but this advantage diminishes in few-shot settings. In the few-shot setting, we observe the “early descent phenomenon” Lin and Lee (2024); Agarwal et al. (2024), where performance initially dips below zero-shot levels with very few examples before improving as the number of shots increases. We ultimately reach 44.4% accuracy at 32-shot. Even at this level, incorporating persona information still yields a further performance gain, suggesting that persona-based and example-based approaches provide complementary signals for modeling individual differences.
@@
The iNews dataset stands to benefit a wide range of research areas: for affective computing researchers modeling emotion recognition while account for individual differences; for LLM developers advancing personalization and subjective phenomena handling; for human behavior simulation researchers modeling individual-level information processing; for social computing scholars investigating demographic effects in content presentation; and for AI alignment researchers studying preference diversity across human populations.
@@ 3 Dataset Collection Protocol
To address the limitations of existing emotion detection datasets, we develop a two-stage data collection protocol to capture individualized affective responses to news headlines (Figure 1). Our protocol emphasizes ecological validity and the collection of rich persona variables to enable the study of how personal characteristics influence affective responses.
@@
Annotators are provided with detailed guidelines, adapted from Bradley and Lang (2007), which are accessible throughout the annotation process [^3] Link: https://docs.google.com/document/d/1RPkjaPSksRbCy3y5d4WltidcUGhlH_np-aAuY2eH33c/. The annotation interface is built using the Potato annotation tool Pei et al. (2022).
We then present annotators with news posts public Facebook pages of major UK outlets (see Table 5). For each news post screenshot (see Figure 8 for exact question wording), annotators provide five types of responses: Dimensional Emotion Ratings capture valence, arousal, and dominance on a Likert scale of 1-7 using the Self-Assessment Manikin (SAM) Bradley and Lang (1994). Discrete Emotion Classifications involve categorizing into one of Plutchik’s eight basic emotion Plutchik (1980). Modality Influence assesses the relative influence of the image versus the text on their emotional response. Personal Relevance rates the headline’s personal relevance. Sharing Likelihood measures the likelihood they would share the post.
@@ 4 Descriptive Analysis
Our dataset comprises 2,899 annotated news posts, with an average of 5.18 annotations per post from 291 distinct annotators. Each annotator also provides ratings for three standardized items from the Affective Norms for English Text (ANET) Bradley and Lang (2007).
@@ 7.1 Experimental Setup
We examine four input conditions. The text-only condition provides a detailed textual description of each news post, while the image-only condition uses the original news screenshot. We then augment each of these base conditions with persona information, creating text-with-persona and image-with-persona conditions where annotator characteristics are incorporated into the system prompt. Since our annotators originally rated news screenshots, we leverage Gemini 1.5 Pro to generate comprehensive textual descriptions for the text-only conditions, enriching these with headline text and engagement metrics. The complete prompt templates and an illustrative news post image-text pair are provided in Sections A.10 and A.9, respectively. We present our zero-shot evaluation results in Table 2.
@@ 8 Conclusion
We introduce iNews, a novel dataset capturing individualized affective responses to news content, demonstrating that annotator characteristics could explain 15.2% of response variance - higher than existing datasets. In our case study, we find that incorporating persona information consistently improves prediction accuracy (up to 7% gains). We observe the early ascent phenomenon, where few-shot performance initially drops below zero-shot levels before recovering with additional examples. Notably, even at 32-shot (achieving 44.4% accuracy), persona information continues to provide benefits, suggesting that explicit modeling of individual characteristics offers complementary signals to behavioral demonstrations. We also identify a modality gap: while image inputs excel in zero-shot settings, they show limited few-shot scaling compared to text inputs. Through its rich annotations and persona variables, iNews advances research in personalization, subjectivity, and affective computing, while providing new opportunities for studying human behavioral simulation.
@@ 9 Limitations
Our emotion measurement framework, while grounded in established psychological constructs (VAD, Plutchik), may not capture the full complexity of emotional responses to news content. Like most studies in this domain, we rely on self-reported emotions, which can be subject to various biases such as social desirability. Future work could strengthen the validity of these measurements by incorporating physiological measures (e.g., skin conductance, facial expressions) or triangulating multiple measurement approaches.
@@ A.3 Geographic Representation of Annotators
A.4 Extended Descriptive Analysis
A.4.1 Additional Annotation Distributions Analysis
@@
A.4.2 Outlet-level Analysis
@@
A.4.3 Relationship Between Arousal and Valence
@@
A.5 Inter-annotator agreement
@@
A.6 Topic Classification Details
@@
A.7 Regression Analysis Details
@@
A.8 Post-Annotation Questionnaire
@@ A.8.7 Methodological Limitation
A.9 Example Image Textual Description
We show an example news post as well as the textual description generated with the prompt in A.6.
@@ A.9 Example Image Textual Description
A.10 Predicting Individual Affective Arousal Prompt
@@
A.11 Additional Few-shot learning results

### v2

@@ Abstract
Understanding how individuals perceive and react to information is fundamental for advancing social and behavioral sciences and developing human-centered AI systems. Current approaches often lack the granular data needed to model these personalized responses, relying instead on aggregated labels that obscure the rich variability driven by individual differences. We introduce iNews, a novel large-scale dataset specifically designed to facilitate the modeling of personalized affective responses to news content. Our dataset comprises annotations from 291 demographically diverse UK participants across 2,899 multimodal Facebook news posts from major UK outlets, with an average of 5.18 annotators per sample. For each post, annotators provide multifaceted labels including valence, arousal, dominance, discrete emotions, content relevance judgments, sharing likelihood, and modality importance ratings. Crucially, we collect comprehensive annotator persona information covering demographics, personality, media trust, and consumption patterns, which explain 15.2% of annotation variance - substantially higher than existing NLP datasets. Incorporating this information yields a 7% accuracy gain in zero-shot prediction and remains beneficial even with 32-shot in-context learning. iNews opens new possibilities for research in LLM personalization, subjectivity, affective computing, and human behavior simulation.
@@
Tiancheng Hu University of Cambridge th656@cam.ac.uk [Uncaptioned image] Data                        Nigel Collier University of Cambridge nhc30@cam.ac.uk [Uncaptioned image] Code
@@ 1 Introduction
Understanding and predicting individual human behavior represents a central challenge across social and behavioral sciences, with applications ranging from public health interventions to economic policy design. Researchers have long recognized that individual characteristics, including personality traits, demographics, cultural backgrounds, and personal experiences, fundamentally shape how people respond to identical stimuli Kring and Gordon (1998); Costa and McCrae (2008); Mesquita and Frijda (1992). However, computational approaches to modeling human responses in natural language processing (NLP) have largely overlooked this rich individual variability, instead relying on aggregated labels that obscure the person-specific patterns that drive real-world behavioral phenomena Plank (2022); Cabitza et al. (2023).
This limitation becomes particularly problematic in NLP applications that aim to understand or predict human responses to textual content. Current approaches typically rely on generic group-level models that ignore individual differences in how people interpret and emotionally respond to identical texts. The emergence of large language models (LLMs), to which psychological theories are increasingly applied for analysis and development Zhu et al. (2024); Liu et al. (2025); Hu et al. (2025), presents new opportunities for modeling these individual-level differences, but requires datasets that capture both behavioral outcomes and the personal characteristics that drive them.
Affective responses to news content represent an ideal testing ground for benchmarking and developing individual-level behavioral simulation models. This domain offers several key advantages: (1) emotional reactions are observable behavioral outcomes with established measurement frameworks Mehrabian and Russell (1974); Bradley and Lang (1994); Ekman (1992), (2) individual differences in affective news processing are well-documented in psychology and media studies Oliver (2002); Valkenburg and Peter (2013); Soroka et al. (2019), (3) the task involves real-world stimuli that people encounter daily. Furthermore, news consumption behavior has direct societal relevance for understanding information processing, media effects, and the development of responsible information systems.
@@
We introduce iNews, a novel large-scale dataset specifically designed to capture the inherent subjectivity of affective responses to real-world news content (overview in Figure 1). Our dataset comprises fine-grained affective responses from 291 annotators to 2,899 Facebook posts from leading UK media outlets. The annotations include: valence-arousal-dominance (VAD) ratings Mehrabian and Russell (1974); Bradley and Lang (1994), Ekman’s basic emotions Ekman (1992), perceived post relevance, modality importance, and sharing likelihood. In addition, we collect a comprehensive set of annotator characteristics [^1] Throughout this work, we also use the term persona information to refer to the same concept. (e.g. demographics, personality, media consumption habits), drawing upon insights from the differential media effects literature.
@@
Our regression analysis confirms that annotator characteristics explain a substantial portion of the annotation variance (15.2% - higher than observed in any NLP dataset to date), highlighting the importance of incorporating individual differences when modeling subjective phenomena like affect. Furthermore, through an open-ended questionnaire with a subset of annotators ($N=20$), we identify nuanced patterns in how individuals experience and articulate their emotional reactions to news content, extending beyond the scope of our structured annotations and survey.
@@
In a case study demonstrating the practical value of this rich persona information, we show that incorporating annotator characteristics can improve LLM predictions of individual-level affective responses by up to 7% in accuracy in zero-shot settings, although overall accuracy remains relatively modest (around 40%). When comparing input modalities (image vs. text), we find that image inputs typically outperform text in zero-shot scenarios but this advantage diminishes in few-shot settings. In the few-shot setting, we observe the “early ascent phenomenon” Lin and Lee (2024); Agarwal et al. (2024), where performance initially dips below zero-shot levels with very few examples before improving as the number of shots increases. We ultimately reach 44.4% accuracy at 32-shot. Even at this level, incorporating persona information yields additional performance gains, suggesting that persona-based and example-based approaches provide complementary signals for modeling individual differences.
@@
The iNews dataset benefits a wide range of research areas: affective computing researchers modeling emotion recognition while accounting for individual differences; LLM developers advancing personalization and subjective phenomena handling; human behavior simulation researchers modeling individual-level information processing; social computing scholars investigating demographic effects in content presentation; and AI alignment researchers studying preference diversity across human populations.
@@ 3 Dataset Collection Protocol
To address the limitations of existing emotion detection datasets and move towards more nuanced individual-level modeling, we develop a two-stage data collection protocol to capture individualized affective responses to news headlines (Figure 1). Our protocol emphasizes ecological validity and the collection of rich persona variables to enable the study of how personal characteristics influence affective responses.
@@
Annotators are provided with detailed guidelines (see Section A.4), adapted from Bradley and Lang (2007), which are accessible throughout the annotation process [^2] Annotators are shown this link: https://docs.google.com/document/d/1RPkjaPSksRbCy3y5d4WltidcUGhlH_np-aAuY2eH33c/. The annotation interface is built using the Potato annotation tool Pei et al. (2022).
We then present annotators with news posts public Facebook pages of major UK outlets (see Table 6). For each news post screenshot (see Figure 8 for exact question wording), annotators provide five types of responses: Dimensional Emotion Ratings capture valence, arousal, and dominance on a Likert scale of 1-7 using the Self-Assessment Manikin (SAM) Bradley and Lang (1994). Discrete Emotion Classifications involve categorizing into one of Ekman’s basic emotion Ekman (1992). Modality Influence assesses the relative influence of the image versus the text on their emotional response. Personal Relevance rates the headline’s personal relevance. Sharing Likelihood measures the likelihood they would share the post. We randomize the order of the news posts.
Quality Control
We recruit our annotators from Prolific, a platform recognized for having high participant attention and comprehension Peer et al. (2022). To further ensure data quality, we implement several procedural checks.
Annotators are required to spend at least 2 minutes on the instruction page (enforced by not showing the continuation button until 2 minutes’ time). The average time spent reviewing the instructions is 4.67 minutes, suggesting reasonable engagement beyond the minimum requirement.
We incorporate a comprehension check at the very beginning of the annotation task. Specifically, we present an excerpt from the annotation manual and ask two questions to see whether the annotators can actually understand our task. Those who fail to answer either one correctly are not allowed to proceed.
We include two attention check questions in our annotation job. Only 7 out of 291 annotators (2.4%) fail either one of these checks, with several contacting us afterward acknowledging their errors. While these annotators are not excluded in the analysis, their IDs are flagged, allowing dataset users to filter them if desired.
To empirically validate that annotators understand and correctly apply the annotation guidelines and the annotation scales, we include three standardized Affective Norms for English Text (ANET) sentences as calibration items Bradley and Lang (2007) in the annotation. Comparing our annotators’ ratings on these items to the original ANET norms (converted to a 7-point scale) provides a direct measure of their ability to use the SAM scales consistently with the established protocol. Due to ANET licensing restrictions, the specific sentences cannot be disclosed. Table 1 presents a comparison of mean scores and standard deviations.
As shown in Table 1, VAD scores from our UK-based annotator pool demonstrate strong correspondence with the original US-based ANET norms for these standardized stimuli. Mean differences are generally minor (typically <0.5 on the 7-point scale). This close alignment indicates that our annotators understand and apply the SAM scales in a manner consistent with established psychometric standards. While minor variations are expected due to demographic differences between our diverse UK sample and the original US university student sample, the overall consistency validates the integrity of our collected affective ratings.
Sentence ID Valence Arousal Dominance
ANET Ours ANET Ours ANET Ours
6400 2.17 (1.24) 1.55 (0.88) 6.32 (1.12) 6.19 (1.05) 2.91 (1.91) 2.57 (1.45)
2550 6.08 (0.98) 6.30 (0.84) 2.15 (1.54) 1.81 (1.24) 5.10 (1.39) 4.38 (0.95)
4450 6.46 (0.83) 6.17 (1.02) 5.91 (1.36) 5.58 (1.14) 4.50 (1.58) 4.91 (1.22)
Table 1: Comparison of Valence (V), Arousal (A), and Dominance (D) ratings between ANET dataset and our annotations. Values show mean scores with standard deviations in parentheses.
@@ 4 Descriptive Analysis
Our dataset comprises 2,899 annotated news posts, with an average of 5.18 annotations per post from 291 distinct annotators.
@@ 7.1 Experimental Setup
We examine four input conditions. The text-only condition provides a detailed textual description of each news post, while the image-only condition uses the original news screenshot. We then augment each of these base conditions with persona information, creating text-with-persona and image-with-persona conditions where annotator characteristics are incorporated into the system prompt. Since our annotators originally rated news screenshots, we leverage Gemini 1.5 Pro to generate comprehensive textual descriptions for the text-only conditions, enriching these with headline text and engagement metrics. The complete prompt templates and an illustrative news post image-text pair are provided in Sections A.11 and A.10, respectively. We present our zero-shot evaluation results in Table 3.
@@ 8 Conclusion
To facilitate the study of individual human behavior in response to stimuli, we introduce iNews, a novel dataset capturing individualized affective responses to news content, demonstrating that annotator characteristics could explain 15.2% of response variance - higher than existing datasets. In our case study, we find that incorporating persona information consistently improves prediction accuracy (up to 7% gains). We observe the early ascent phenomenon, where few-shot performance initially drops below zero-shot levels before recovering with additional examples. Notably, even at 32-shot (achieving 44.4% accuracy), persona information continues to provide benefits, suggesting that explicit modeling of individual characteristics offers complementary signals to behavioral demonstrations. We also identify a modality gap: while image inputs excel in zero-shot settings, they show limited few-shot scaling compared to text inputs. Through its rich annotations and persona variables, iNews advances research in personalization, subjectivity, and affective computing, while providing new opportunities for studying human behavioral simulation.
@@ 9 Limitations
Our emotion measurement framework, while grounded in established psychological constructs (VAD, Ekman), may not capture the full complexity of emotional responses to news content. Like most studies in this domain, we rely on self-reported emotions, which can be subject to various biases such as social desirability. Future work could strengthen the validity of these measurements by incorporating physiological measures (e.g., skin conductance, facial expressions) or triangulating multiple measurement approaches. Additionally, it is important to acknowledge that this paper is concerned with affective responses to news, which may differ in nature or intensity from emotions stemming from more direct, grounded lived experiences.
@@ A.3 Geographic Representation of Annotators
A.4 Annotation Manual
A.5 Extended Descriptive Analysis
A.5.1 Additional Annotation Distributions Analysis
@@
A.5.2 Outlet-level Analysis
@@
A.5.3 Relationship Between Arousal and Valence
@@
A.6 Inter-annotator agreement
@@
A.7 Topic Classification Details
@@
A.8 Regression Analysis Details
@@
A.9 Post-Annotation Questionnaire
@@ A.9.7 Methodological Limitation
A.10 Example Image Textual Description
We show an example news post as well as the textual description generated with the prompt in A.7.
@@ A.10 Example Image Textual Description
A.11 Predicting Individual Affective Arousal Prompt
@@
A.12 Additional Few-shot learning results